Let  Select all that applyLet z =f(x,y)= arctan(3x In(6) Select all that apply Your answer: The slope of the tangent line to the curve obtained by inte
Answer #1

o fix,y)= tant/34/n(64) fo(a,y)= 3 in (64) 1492?l?6) x 6X2x1n (6) fryczy)= (1+9a\u2@y) gr 3 en 6y) by (1 +92 m (6) 3 - 272m (kindly please upvote this

Option c, e and f are correct.